### v1

# RATSF: Empowering Customer Service Volume Management through Retrieval-Augmented Time-Series Forecasting

Abstract
An efficient customer service management system hinges on precise forecasting of service volume. In this scenario, where data non-stationarity is pronounced, successful forecasting heavily relies on identifying and leveraging similar historical data rather than merely summarizing periodic patterns. Existing models based on RNN or Transformer architectures often struggle with this flexible and effective utilization. To address this challenge, we propose an efficient and adaptable cross-attention module termed RACA, which effectively leverages historical segments in forecasting task, and we devised a precise representation scheme for querying historical sequences, coupled with the design of a knowledge repository. These critical components collectively form our Retrieval-Augmented Temporal Sequence Forecasting framework (RATSF). RATSF not only significantly enhances performance in the context of Fliggy hotel service volume forecasting but, more crucially, can be seamlessly integrated into other Transformer-based time-series forecasting models across various application scenarios. Extensive experimentation has validated the effectiveness and generalizability of this system design across multiple diverse contexts.
1 Introduction
At the heart of customer service volume management lies the accurate estimation of total service demand, which significantly impacts system costs. For every underestimation of 100 service requests, there is a need to urgently mobilize and supplement service staff at a cost equivalent to three times the single-unit labor cost; conversely, overestimating by 100 requests results in a waste of that same single-unit labor cost. Service volume forecasting for the Travel industry is particularly intricate because of its inherent interconnections with numerous variables, including order statuses, weather variations, international developments, and destination country policies. These factors contribute to highly non-stationary data patterns that, in turn, can lead to considerable biases when employing traditional uni-variate time series forecasting methods [9, 1, 23] which rely on period summaries and trend analyses.
In fact, across domains like stock market analysis and station traffic forecasting where business cycles and fluctuations play a significant role, there is a shared desire for a universal and flexible time series forecasting system design that can adeptly utilize historical sequence information to tackle intricate prediction challenges. Meanwhile, in the realm of time series forecasting(TSF), where models, whether based on Recurrent Neural Networks (RNNs) or Transformer, commonly face difficulties efficiently processing and extracting insights from vast amounts of historical data.
One naive way to solve this is to try elongating the sequence that the transformer processes. Some approach involves sampling historical data to fit within a limited context window. Specifically, the Informer [32] algorithm samples K points from the sequence and derives a shorter Q sequence based on these sampled points. However, this approach assumes that all historical information is equally important, which may not be suitable for many time series scenarios where different data points can carry varying significance. Perceiver [10] and similar methods opt for a different approach by mapping Query sequences to fixed lengths, reducing computation and allowing for more historical data storage. Nonetheless, they still face challenges in efficiently extracting and interpreting critical information from extended time series.
In the field of natural language processing, strides have been made to expand Transformer models’ context handling. On one front, techniques like flash-attention [5] enhance efficiency and reduce complexity, enabling longer context processing. However, it’s crucial to note that directly extending receptive fields may cause larger models to overlook significant details in lengthy inputs, like [15] mentioned.
On another front, Retrieval-Augmented Generation (RAG)-like methods have drawn attention for enhancing model performance by incorporating external information. NVIDIA’s [2] research indicates that even when models already handle large context windows in text tasks, they can still achieve substantial performance gains by retrieving and using relevant data from external sources.
Employing the concept of RAG, we have identified a potential approach. We can build a knowledge base that encompasses all historical time series, and pinpoint and utilize the most predictive historical segments with an improved transformer model, thereby refining forecast accuracy. To accomplish these, we focused on implementing three central enhancements: a knowledge base schema for efficiently indexing historical data, a cross-attention module to integrate historical information, and a predictive embedding for retrieval purposes.
Figure 1: RATSF uses a TSKB to store and index historical sequences, alongside a transformer-based forecasting model. As illustrated in the bottom-right corner, TSKB segments the full history for efficient retrieval. To do forecasting, the Encoder fetches $\mathbf{X}_{o}$ from t recent time steps, and forms a retrieval sequence with d latest steps, retrieves N related sequences $\mathbf{X}_{r}$ from the TSKB. RACA in the Decoder then merges $\mathbf{X}_{r}$ and $\mathbf{X}_{o}$ info to deliver result.
In concise terms, our main contributions are:
1. We offer a general retrieval-augmented time-series forecasting strategy, RATSF, that can be implemented across a variety of Transformer models and their derivatives to enhance performance.
2. We employ a novel embedding strategy for precise historical piece retrieving.
3. We present a straightforward and manageable design for a time-series knowledge base, which significantly facilitates efficient management of historical data.
4. The method presented in this work currently serves as the operational algorithm for Fliggy’s service volume management system across all business domains, supporting not only hotel online bookings but also support businesses such as visa processing, flight and train ticket reservations, as well as entrance tickets to resort areas. Our method has demonstrated exceptional effectiveness in various practical application scenarios.
2 Review
Despite intense competition [20, 6, 30, 8], Transformer models [25] and their improved variants [19, 26, 16, 13, 14] have become the mainstream choice in time-series forecasting tasks. However, the computational complexity of the original Transformer model scales as O(n²) with respect to sequence length, significantly limiting the maximum sequence length it can handle and thereby constraining its ability to use historical data.
Improvements in Time-Series Forecasting with Transformers. Numbers of works focus on enhancing the Transformer model’s capability to extract temporal features, thereby improving prediction accuracy, exemplified by Fedformer’s [33] introduction of a frequency-augmented Attention mechanism that directly incorporates Fourier operators, a similar path taken by TimesNet [27]. In contrast, Autoformer [28] proposes an operator that decomposes time series information into trend and periodic components. These two types of solutions perform well for relatively stationary sequences but may see performance drop when dealing with highly non-stationary time series data.
Another core strategy involves learning a robust representation of the time series first, which is then used to enhance the forecasting accuracy. TNC [24] harnesses the core concept of contrastive learning and employs samples within a specific temporal neighborhood within a window as positive pairs, while treating samples from differing temporal neighborhoods as negative pairs. Despite these improvements having collectively boosted the ability of Transformers in time-series forecasting tasks, none has fundamentally increased the Transformer’s receptive field.
Retrieve Augmented Method in NLP. In last two years, the Retrieval-Augmented Generation (RAG) approach [7, 4, 18, 31, 21, 12, 34] has gained widespread adoption in the field of NLP. REALM harnesses a knowledge retriever to distill information from vast corpora and thereby enhance the performance of pre-trained language models. Transformer-XL+kNN [3] incorporates a K-Nearest Neighbors algorithm to search through training data, refining dialogue generation capabilities. In the context of named entity recognition tasks, U-RaNER [22] utilizes multi-modal heterogeneous retrieval techniques to boost knowledge retrive and, by integrating retrieved knowledge into the model, strengthens its understanding of queries and improves entity recognition accuracy.
Retrieval Augmented Method in Time-series Forecasting. MQ-ReTCNN [29] is designed for complex time series prediction tasks involving multiple entities and variables. It employs a scoring function to compile relevant contexts from offline data, selects scored segments, and appends them to the prediction sequence.Its retrieval mechanism emphasizes leveraging historical sequences of one entity to enhance predictions about another, without directly utilizing historical data for direct prediction assistance.
ReTime [11] creates a relation graph based on temporal closeness between sequences and employs relational retrieval instead of content-based retrieval.It does not optimally use historically similar sequences as reference points due to its inherent design limitations. Both MQ-ReTCNN and ReTime incorporate retrieval enhancement strategies but have yet to introduce a general and efficient retrieval technique specifically for single-variable time series prediction scenarios.
Figure 2: Each RACA has two cross-attention modules: one utilizes the Encoder output as K,V, while the other employs embedded retrieved sequences as K,V. Both outputs are concatenated and passed through a Linear module to reshape back to the input dimensions.
3 Method
Subsequently, we first provide an overview of the RATSF’s overall structure and then examine the key features of each component in detail.
The simplest description of RATSF is to first build a TSKB and then retrieve relevant sequences from it to enhance the forecasting model. For a Time Series Forecasting (TSF) task within a specific context, we first construct a TSKB using the schema detailed in Section 3.2. Next, we utilize a retrieval method to gather relevant sequences, with this method being introduced in Section 3.3. Finally, we integrate these retrieved sequences into the forecasting model, with its core computational component, RACA, being introduced in Section 3.1. Figure 1 shows the data flow of RATSF.
3.1 RACA
As previously mentioned, Retrieval-Augmented Cross-Attention (RACA) is a module designed to be plugged-into a time-series forecasting model. In our demonstration, we employ the vanilla Transformer as the time-series forecasting model; however, it’s important to note that the RACA module is designed to be compatible and can be seamlessly plug into any transformer-based time-series model.
we now provide an explanation of the abbreviations represented by subscript and superscript symbols in the following paragraphs: The letter ’o’ for original, represents the input time-series piece entering the encoder; letter ’b’ for embedding , letter ’e’ for encoder; letter ’r’ for retrieved sequence, letter ’f’ for sequence for forecasting; ’l’ for layer ; ’D’ for hidden dimension.
In transformer architecture, an encoder consists of an embedding module and multiple self-attention modules. Initially, the raw time series data $\mathbf{X}_{o}=\left\{x_{1},...,x_{t}\right\}$, $\mathbf{X}_{o}$ when processed through the embedding module, generates an output as the $\mathbf{X}^{b}_{o}$; subsequently, the output resulting from all the encoder calculation is referred as the $\mathbf{X}^{e}_{o}$.
RACA is introduced within the decoder part: besides conventionally utilizing the encoder output and the place holding embedding for the forecasting sequence, it incorporates information from N most relevant history sequences retrieved from TSKB. Like shown in Figure 2, these retrieved sequences are transformed through the encoder’s embedding module, generating Ns $\mathbf{X}^{e}_{r}$  [^1] In section 4, when comparing four Transformer models horizontally, the impact of the Encoder extends to both representation retrieval and decoding stages. To isolate the contribution of the Encoder to the retrieval performance, we employ the Encoder output as the retrieval scheme while incorporating the representation $\mathbf{X}^{b}_{r}$ rather than $\mathbf{X}^{e}_{r}$ into RACA. In section 3.3, we directly compare the performance of using DTW alone versus Encoder output for retrieval. When integrated into the Decoder, the Encoder output’s affecting both retrieval effectiveness and decoding quality. Given the significant differences between Encoders trained solely with DTW and those following our proposed method,as shown in B.2. To ensure consistency across experiments, $\mathbf{X}^{e}_{r}$ is consistently replaced with $\mathbf{X}^{b}_{r}$ in this paper’s all experiments., and we concatenate these N vectors into one with shape$[1,N*L_{r},D]$ for later use.
In Transformer-based time series forecasting models, placeholders are embedded into a $[1,L_{f},D]$ matrix, $\mathbf{H}_{f}$. The decoder consists of $l$ stacked RACA modules for inputs of length $L_{f}$. Each module has two parallel units: Unit 1, like in Function (1), uses $\mathbf{H}_{f}$ as Query with Key and Value from $\mathbf{X}^{e}_{o}$, outputting $\mathbf{H}^{l,1}_{f}$ of the same shape. Unit 2, retrieval-augment part, like in Function (2) ,also queries $\mathbf{H}_{f}$ but fetches Keys and Values from the concatenated sequence of $\mathbf{X}^{b}_{r}$, generating $\mathbf{H}^{l,2}_{f}$. $s$ in the Functions is the scaling factor.
$\mathbf{H}^{l,1}_{f}=\text{softmax}\left(\frac{\mathbf{H}_{f}\mathbf{X}^{e}_{o}}{s}\right)\mathbf{X}^{e}_{o},$ (1)
$\mathbf{H}^{l,2}_{f}=\text{softmax}\left(\frac{\mathbf{H}_{f}\mathbf{X}^{e}_{r}}{s}\right)\mathbf{X}^{e}_{r},$ (2)
$\mathbf{H}^{l+1}_{f}=\delta\left(\text{concat}\left(\mathbf{H}^{l,1}_{f},\mathbf{H}^{l,2}_{f}\right)\right).$ (3)
As shown in Function 3, after concatenating both intermediate vectors,$\mathbf{H}^{l,1}_{f}$ and $\mathbf{H}^{l,2}_{f}$, along the sequence dimension, forms a $[1,2*L_{f},D]$ vector, it undergoes linear transformation before being passed to the next layer.
3.2 TSKB
In the development of TSKB, we employ an innovative strategy for constructing the indexing sequence (K) and content sequence (V). Unlike traditional methods that solely store and retrieve the entire original sequence, TSKB preserves the original sequence as the core content while selects certain segments from it to construct a discriminative indexing sequence. This approach can be likened to processing a written piece where conventional methods involve direct full-text searches for desired information, which may be inefficient and less precise; whereas with TSKB, it’s akin to extracting key headlines from the body of the text to serve as indices that facilitate rapid access to core information. This dual-sequence structure allows the system, when handling large-scale data, to efficiently locate and access targeted portions of the original sequence through customized index sequences, thereby significantly enhancing overall performance.
In this article, to illustrate our system design, we employ a fixed slicing logic to select segments from V for computing the retrieval representation. In section 3.3, we will present in detail how the representation of V is learned.
Figure 3: Within the cold start stage, we use raw sequence of retrieval input to find Top-K relevant sequences using DTW. After one epoch, we switch to Euclidean Distance for retrieval embedding to gather Top-K matches.
Figure 4: TSKB utilizes a rolling window of length $L_{v}$ to collect V, with an indexing segment of length $L_{r}$ taken from its leading part as K , and advances the window in steps of size S.
3.3 Learning Retrieval Embedding
Judging the quality of a retrieval representation mainly depends on how well it captures similarities at key information points relevant to forecasting tasks. Time series data is complex due to numerous information points and the challenge of presetting comparative weights. Therefore, we use this principle: if a representation closely mirrors the important details needed for prediction, its overall performance will be better. Later, we’ll explain the exact steps to put this principle into practice. We acquire the retrieval embedding for TSKB through two procedures.
3.3.1 Sequential Slicing
To obtain content sequence V, we utilize a rolling window approach where the step size is $S$ and the window length is $L_{v}$, to sample all historical sequences and incorporate them into TSKB. Since prediction models only use present data to infer the future, we align indexing sequence K with this constraint. We extract the initial segment of each V, having a length of $L_{r}$, to serve as its K. The selection of $L_{v}$, $L_{r}$ and $S$ is introduced in section 4.
3.3.2 Embedding Learning
To guarantee the embedding in capturing the essential information for forecasting, we utilize the encoder of the RATSF’s forecasting model. The model inherently excels at selecting precise information during training to improve forecast accuracy. In practice, every indexing sequence K from the knowledge base undergoes the RATSF forecasting model’s encoder layer to generate a retrieve embedding.
Meanwhile, in the early stages of training the forecasting model, its encoder is not yet capable of producing representations that accurately match target sequences. The effectiveness and learning pace of the RATSF’s forecasting model are significantly correlated with whether or not it retrieves historical sequences that are beneficial to the forecasting task.
To avoid time and data-consuming iterations stemming from a random initialization state, we introduce DTW as an auxiliary tool during the initial phase of model training, as shown in left part of Figure 3. Initially, DTW is utilized for similarity-based sequence retrieval, which is then incorporated into the iterative training process of the RATSF forecasting model. After finish one epoch of training, we transition to using embedding generated by the forecasting model itself for similar sequence retrieval, continuing the training until the model converges.
Figure 5: The lower plot displays the top 3 relevant retrieved sequences for one sample that are end-to-end connected. Each retrieved sequence has a length of 21 ($L_{f}$), comprising 14 points preceding the forecasting date and 7 points to be forecasted, highlighted with an orange background. The color bars represent aggregated cross-attention values at the first forecasting point, in which lighter color indicate higher attention value. In the upper part of the figure, predicted (yellow) and ground truth (orange) values align at corresponding positions with retrieved sequence. For this sample, the model notably focuses on segments indicating ascent and predicts a subsequent increase in value.
Table 1: Performance comparison of four Transformer models with and without the application of RATSF on three experimental datasets. All results are evaluated using MSE (Mean Squared Error) and MAE (Mean Absolute Error), and the best result is bolded.
models transformer nstransformer autoformer fedformer
metric mse mae mse mae mse mae mse mae
FHSV baseline 4344450.512 1349.146 4414326.000 1242.128 6096673.500 1752.798 5517485.000 1638.275
RATSF 3421546.027 1101.172 3823635.000 1070.933 5597297.500 1657.561 5386818.500 1595.657
ETTh1 baseline 3.187 1.391 1.501 0.920 8.651 2.264 4.159 1.597
RATSF 2.366 1.180 1.407 0.886 7.300 2.113 3.952 1.565
Exchange baseline 0.00165 0.03134 0.00014 0.00922 0.00058 0.01906 0.00172 0.03493
RATSF 0.00068 0.02012 0.00013 0.00912 0.00067 0.02000 0.00068 0.02084
4 Experiments
In order to evaluate the generalizability and efficacy of RATSF in industry applications, we initially carry out a range of experiments to contrast the performance changes among typical time-series models when augmented with RATSF across various tasks, focusing particularly on the Fliggy Hotel Service Volume Dataset (FHSV). Following this, we employ a sequence of ablation studies to separately demonstrate the design choice of each improvement – RACA, TSKB, and Retrieval Embedding.
4.1 Experiment Setup
Datasets. Given that our primary focus lies in the forecasting task within the context of customer service volume management, we initially employed the Fliggy Hotel Service Volume Dataset to substantiate the efficacy of RATSF, the detail description of the dataset is in Appendix A. Additionally, to demonstrate the performance of our approach in other contexts, we employed two classic datasets: ETT and Exchange.
Models.We contrasted the vanilla Transformer against three of its prominent time-series domain variants: Fedformer, Autoformer, and NS-Transformer. Specifically, in keeping with the intrinsic sequential decomposition characteristics of Fedformer and Autoformer, historical segments were initially broken down using each model’s native calculations, for example the seasonal init and trend-cyclical init in Autoformer. Subsequently, these decomposed segments were embedded and fed into the RACA module for further processing.
Forecasting Setting.Consistent with the original intent of validating the effectiveness of RATSF in the context of customer service volume management, we adopt the business setting from Fliggy. In the Fliggy’s practical operations, we are required to finalize our forecasting for the upcoming week’s day-by-day staffing needs one week in advance, which is why the actual prediction window period is set at 7 days, represented as 7 tokens. In the travel industry, which Fliggy App’s serving, a quarter generally defines a relatively complete business trend. Therefore, we choose an encoder window length of 98 days—exceeding 90 days and divisible by 14—which consequently results in $\mathbf{X_{o}}$ is 98 units long. And we concate $\left\{x_{t-14},...,x_{t}\right\}$ with a placeholder of length 7 to form $\mathbf{X_{f}}$. To match the length of $\mathbf{X_{f}}$, we set $L_{v}$=21 and $L_{r}$=14. Regarding normalization, we employ unified $\mu$ and $\sigma$ values for whitening operations during the preprocessing stage and inverse normalization is carried out after forecasting. We have applied the same data processing scheme to both the ETT and Exchange datasets.
Training Setting.We have selected the Adam optimizer, with the batch size$=$64 and max training epochs$=$10. The initial learning rate (lr) is set to 0.0001, employing a linear decay strategy where the decay parameter for the first 5 epochs is 0.9, for the last 5 epochs is 0.5. We also incorporate $L1$ regularization with $\lambda$=0.0001. Concurrently, an early stopping mechanism is activated when the loss ceases to decrease.
Metric.To ensure that our evaluation metrics directly reflect business performance, we first inverse normalize the forecast result as above mentioned and then compute the MSE (Mean Squared Error) and MAE (Mean Absolute Error) against the Ground Truth. The lower these two indicators are, the more accurate the prediction results prove to be. Moreover, each decrement of 1 unit in MAE signifies a corresponding reduction of 1 unit in service staff management expenditure. This configuration has been applied uniformly across all experiments.
4.2 Main Restult
Table 1 presents the performance comparison of four Transformer models with and without the application of RATSF on three experimental datasets. In the scenario of forecasting with the Fliggy Hotel Customer Service Volume Dataset, the Transformer model’s MAE loss dropped by 18% after incorporating RATSF, potentially translating to a reduction of approximately 200 redundant personnel cost in practical management operations. Notably, even with the NSTransformer [17], Autoformer, and FedFormer models, which have undergone temporal optimization in their attention operators, adopted the RATSF design, they respectively recorded decreases in MAE loss by 14%, 5%, and 4%.
Furthermore, cross-comparison reveals that Autoformer and FedFormer yield slightly inferior results compared to the Transformer, mainly due to the stronger non-stationarity inherent in customer service volumes data within the hotel industry. As an illustrative example, while Mondays typically exhibit similar cyclical characteristics, actual service volume during Mondays preceding a short holiday can surge significantly compared to regular ones. In such instances, historical data from periods close to other short holidays provide more valuable insights than simple temporal periodicity and short-term trends; specific case studies will be showcased later.
Additionally, across the ETT and Exchange datasets, the adoption of RATSF design in the Transformer and its time-series optimized variants led to noticeable reductions in MSE metrics. Moreover, among these, the NSTransformer consistently outperforms the other models, likely because its mean-adapted feature proves effective in a broad range of time-series prediction domains.
4.3 Ablation Study
Section 4.3.1 comparatively analyzes the differences in various designs for integrating historical pieces into forecasting. Section 4.3.2 examines the impact of integrating varying numbers of historical pieces on forecast effectiveness. Section 4.3.3 demonstrates the advantages of utilizing model encoder for recalling historical pieces.
Table 2: Ablation of the specifically designed structures within RACA on FHSV scenario.
dataset methods metric
mse 4344450.512
baseline mae 1349.146
baseline + mse 4071384.75
enhanced encoder mae 1226.850
mse 3421546.027
FHSV baseline + RACA mae 1101.172
4.3.1 RACA’s design choice
Based on the Transformer architecture, two designs are considered for integrating historical pieces into the forecasting model: Design One represents a straightforward approach where historical pieces are combined with the forecasting context inputs within the encoder; Design Two employs our advanced RACA design that integrate historical pieces using cross-attention mechanisms within the decoder.
With the Fliggy Hotel Service Volume Dataset, we contrasted the performance of three methods: No historical pieces used(baseline experiment), Strategy One, and Strategy Two (RACA). The outcomes in Table 2 demonstrate that the RACA design excels, with its MAE notably lower than both the baseline and Strategy One.
A deeper examination disclosed that the Cross Attention mechanism in the RACA design efficiently guides the forecast sequence to concentrate on past trends analogous to the current scenario. Figure 5 illustrates the attention map within the RACA for the data points under forecast, where highlighted color blocks indicate that the data point being forecast places significant attention on historical segments that exhibiting similar trend patterns, and these historical segments’ trends closely align with the actual waveform following the data point in the ground truth.
4.3.2 Optimal Retrieval Count
Table 3: Results of recalling different numbers of pieces in our RATSF model.
dataset methods metric
mse 4344450.512
FHSV baseline mae 1349.146
mse 3678230.020
recalled Top-1 pieces mae 1160.101
mse 3700934.754
recalled Top-2 pieces mae 1155.952
mse 3421546.027
recalled Top-3 pieces mae 1101.172
mse 3460814.538
recalled Top-4 pieces mae 1108.895
mse 3750426.751
recalled Top-5 pieces mae 1231.564
Having established that integrating effective historical pieces positively impacts forecasting, we then inquire about the optimal number of sequences to retrieve. To demonstrate the effect of this choice, we gradually increase the number of integrated retrieve sequences from 0 to 5. As shown in Table 3, during the process of expanding from integrating no historical piece to integrating up to three historical pieces, the MAE error consistently decreases; however, when the quantity of integrated historical pieces is further increased, the forecasting error starts to escalate.
Upon analyzing samples, the insight is obvious: not all of the top five retrieved historical piece for most of samples closely resemble the Ground Truth. The reason being, focusing solely on the TopK ranking without adequately considering similarity thresholds can easily lead to a scenario where, as K increases, the quality of the retrieved pieces becomes less assured and more prone to introduce noise and irrelevant information, thus potentially causing confusion instead.
4.3.3 Retrieval Embedding
We claimed that DTW is not optimal as the ultimate retrieving method. In this part of ablation we show that with an experiment on FHSV. We compare DTW, a trained MLP and current design choice, the forecasting model’s encoder in terms of accuracy.
Table 4 clearly demonstrates that the forecasting model’s encoder output provide embedding outperforms both the other two schemes significantly. The advantages of encoder output over MLP are evident, particularly because the attention mechanism within the transformer’s encoder is better suited to discern and refine the interdependence’s among data points in sequences, thereby abstracting more accurate patterns.
Table 4: Comparison of several variants of retrieval representation schemes.
dataset methods metric
mse 4344450.512
baseline mae 1349.146
mse 3787191.250
baseline + DTW mae 1226.690
mse 3784188.541
baseline + MLP mae 1210.785
mse 3421546.027
FHSV baseline + encoder mae 1101.172
5 Conclusion
In this paper, we propose the RATSF and explain its effectiveness in enhancing time-series forecasting performance. Specifically, our approach improves from three dimensions: first, we integrate the RACA mechanism into the transformer model architecture, which leverages historical data segments similar to current contexts to bolster forecasting prowess. Second, we construct a KV schema for a time-series knowledge base that significantly enhances the flexibility and precision of the retrieval process. And third, we utilize the Encoder component within the forecasting model to embed K sequences, thereby refining the accuracy of retrieving relevant historical information. The method introduced in this study has been actively operational within Fliggy’s service volume management system for several months. Of particular importance is that due to the strong flexibility and scalability inherent in this design, the method can play a significant role across a wide array of practical application scenarios.
Appendix A Fliggy Hotel Service Volume Dataset
The Fliggy Hotel Service Volume Dataset (FHSV) is derived from service records within the Fliggy APP’s customer service center, capturing the volume of daily customer service operations related to inquiries about hotel reservations, cancellations, and changes. Each instance of a service request received either through the APP interface or by phone is counted as a single service event. Daily aggregates of such events are recorded, forming individual data points. This dataset comprises 1740 data points collected between January 1, 2019, and October 7, 2023.
To facilitate experimentation, this time series dataset has been sequentially partitioned in a chronological order into three subsets: the train set (from January 1, 2019, up until May 24, 2022), the evaluation set (from May 25, 2022, to October 31, 2022), and the test set (from November 1, 2022, onward). The reported experimental results are based on the model’s performance on the test set alone.
Appendix B Other Experiment Details
B.1 Length of Retrieval Index Segment of Knowledge Base
Table 5: Length of Retrieval Index Segment of Knowledge Base.
dataset retrieval length metric
mse 4621934.067
4 mae 1303.329
mse 3828479.700
7 mae 1178.574
mse 3421546.027
14 mae 1101.172
mse 3747974.421
21 mae 1208.379
mse 3816283.253
28 mae 1243.133
mse 3913793.800
FHSV 35 mae 1332.472
The length of the retrieval segment in a knowledge base significantly affects the accuracy of retrieving the original sequence. Through controlled experiments, we identified the most suitable retrieval segment length for the Fliggy scenario. As shown in Table 5, we incrementally extended the length of the retrieval segment K from 4 units to 35 (corresponding to half to five times the length of the prediction sequence at 7 units), then observed the impact on model performance as measured by MSE and MAE.
The results revealed that during the expansion of the retrieval segment from 4 to 14 units, the predictive performance improved. However, beyond this point, as the retrieval segment continued to lengthen, the predictive performance began to decline. This trend demonstrates that, in practical scenarios, as the retrieval segment grows from short to long, its representation transitions from being information-poor to increasingly mixed and complex.
This situation underscores why, in traditional time series databases, using full-ordered representations for retrieval often yields effective outcomes. By contrast, our K,V design effectively circumvents this issue, providing flexible and accurate retrieval results.
B.2 Effectiveness of Using DTW as Auxiliary Training
In the article, we claimed the effectiveness of using DTW as a feature retrieval technique during the cold start phase to aid in single-epoch training for forecasting models. To validate the necessity of this approach, we designed a set of comparative experiments: a baseline group without DTW-assisted training (named by end to end), and other groups that were each trained with DTW assistance for 1 to 5 epochs. We ensure all experiments had same maximum numbers of training epochs and an identical early stopping strategy. Table 6 shows that the model’s predictive performance is poorest when DTW is not employed; moreover, no better results were obtained when DTW-assisted training exceeded one epoch.
This finding confirms our expectation that, without DTW support during the initial stage, the model struggles to discover high-quality retrieval vectors. Concurrently, our main experiment has previously demonstrated that DTW is not inherently an optimal retrieval scheme geared towards forecasting accuracy. Consequently, prolonging the learning period under DTW guidance can actually restrict the expressive capacity of the forecasting model rather than improving it.
Table 6: Effectiveness of Using DTW as Auxiliary Training.
dataset number of epochs metric
mse 3421546.027
1 epoch mae 1101.172
mse 3769362.251
2 epoch mae 1252.136
mse 3697616.753
3 epoch mae 1135.631
mse 3720408.244
4 epoch mae 1248.433
mse 3842794.029
5 epoch mae 1278.324
mse 3800124.170
FHSV end to end mae 1302.054

### v2

@@
 RATSF: Empowering Customer Service Volume Management through Retrieval-Augmented Time-Series Forecasting
-Abstract
+Abstract.
 An efficient customer service management system hinges on precise forecasting of service volume.
 In this scenario, where data non-stationarity is pronounced, successful forecasting heavily relies on identifying and leveraging similar historical data rather than merely summarizing periodic patterns.
-Existing models based on RNN or Transformer architectures often struggle with this flexible and effective utilization.
-To address this challenge, we propose an efficient and adaptable cross-attention module termed RACA, which effectively leverages historical segments in forecasting task, and we devised a precise representation scheme for querying historical sequences, coupled with the design of a knowledge repository.
-These critical components collectively form our Retrieval-Augmented Temporal Sequence Forecasting framework (RATSF).
-RATSF not only significantly enhances performance in the context of Fliggy hotel service volume forecasting but, more crucially, can be seamlessly integrated into other Transformer-based time-series forecasting models across various application scenarios.
+Existing models based on RNN or Transformer architectures may struggle with this flexible and effective utilization.
+To tackle this challenge, we initially developed the Time Series Knowledge Base (TSKB) with an advanced indexing system for efficient historical data retrieval.
+We also developed the Retrieval Augmented Cross-Attention (RACA) module, a variant of the cross-attention mechanism within Transformer’s decoder layers, designed to be seamlessly integrated into the vanilla Transformer architecture to assimilate key historical data segments.
+The synergy between TSKB and RACA forms the backbone of our Retrieval-Augmented Time Series Forecasting (RATSF) framework.
+Based on the above two components, RATSF not only significantly enhances performance in the context of Fliggy hotel service volume forecasting but also adapts flexibly to various scenarios and integrates with a multitude of Transformer variants for time-series forecasting.
 Extensive experimentation has validated the effectiveness and generalizability of this system design across multiple diverse contexts.
-1 Introduction
-At the heart of customer service volume management lies the accurate estimation of total service demand, which significantly impacts system costs.
-For every underestimation of 100 service requests, there is a need to urgently mobilize and supplement service staff at a cost equivalent to three times the single-unit labor cost; conversely, overestimating by 100 requests results in a waste of that same single-unit labor cost.
-Service volume forecasting for the Travel industry is particularly intricate because of its inherent interconnections with numerous variables, including order statuses, weather variations, international developments, and destination country policies.
-These factors contribute to highly non-stationary data patterns that, in turn, can lead to considerable biases when employing traditional uni-variate time series forecasting methods [9, 1, 23] which rely on period summaries and trend analyses.
+PVLDB Reference Format:
+PVLDB, 14(1): XXX-XXX, 2020.
+doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License.
+Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
+For any use beyond those covered by this license, obtain permission by emailing info@vldb.org.
+Copyright is held by the owner/author(s).
+Publication rights licensed to the VLDB Endowment.
+Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
+doi:XX.XX/XXX.XX
+1. Introduction
+Accurate estimation of total service demand is crucial for customer service volume management in the travel industry, including hotels, airlines, attractions, and transaction-brokering apps like Fliggy, significantly affecting system costs.Underestimating by 100 service requests incurs urgent mobilization costs, equivalent to three times the labor cost for a single request; overestimation, meanwhile, leads to wasted labor costs.
+The travel industry presents unique challenges in service volume forecasting due to its interplay with variables like order statuses, weather, international events, and destination country policies.
+At Fliggy, precise forecasting is essential for planning the recruitment and training of customer service staff, as well as scheduling.
+These factors result in non-stationary data patterns that can introduce significant biases with traditional uni-variant time-series forecasting methods (Hyndman and Athanasopoulos, 2018; Anderson and Kendall, 1976; Taylor and Letham, 2018), which rely on periodic summaries and trend analyses.
 In fact, across domains like stock market analysis and station traffic forecasting where business cycles and fluctuations play a significant role, there is a shared desire for a universal and flexible time series forecasting system design that can adeptly utilize historical sequence information to tackle intricate prediction challenges.
@@
 Some approach involves sampling historical data to fit within a limited context window.
-Specifically, the Informer [32] algorithm samples K points from the sequence and derives a shorter Q sequence based on these sampled points.
+Specifically, the Informer (Zhou et al., 2021) algorithm samples K points from the sequence and derives a shorter Q sequence based on these sampled points.
 However, this approach assumes that all historical information is equally important, which may not be suitable for many time series scenarios where different data points can carry varying significance.
-Perceiver [10] and similar methods opt for a different approach by mapping Query sequences to fixed lengths, reducing computation and allowing for more historical data storage.
+Perceiver (Jaegle et al., 2021) and similar methods opt for a different approach by mapping Query sequences to fixed lengths, reducing computation and allowing for more historical data storage.
 Nonetheless, they still face challenges in efficiently extracting and interpreting critical information from extended time series.
 In the field of natural language processing, strides have been made to expand Transformer models’ context handling.
-On one front, techniques like flash-attention [5] enhance efficiency and reduce complexity, enabling longer context processing.
-However, it’s crucial to note that directly extending receptive fields may cause larger models to overlook significant details in lengthy inputs, like [15] mentioned.
+On one front, techniques like flash-attention (Dao et al., 2022) enhance efficiency and reduce complexity, enabling longer context processing.
+However, it’s crucial to note that directly extending receptive fields may cause larger models to overlook significant details in lengthy inputs, like (Liu et al., 2023) mentioned.
 On another front, Retrieval-Augmented Generation (RAG)-like methods have drawn attention for enhancing model performance by incorporating external information.
-NVIDIA’s [2] research indicates that even when models already handle large context windows in text tasks, they can still achieve substantial performance gains by retrieving and using relevant data from external sources.
-Employing the concept of RAG, we have identified a potential approach.
-We can build a knowledge base that encompasses all historical time series, and pinpoint and utilize the most predictive historical segments with an improved transformer model, thereby refining forecast accuracy.
-To accomplish these, we focused on implementing three central enhancements: a knowledge base schema for efficiently indexing historical data, a cross-attention module to integrate historical information, and a predictive embedding for retrieval purposes.
-Figure 1: RATSF uses a TSKB to store and index historical sequences, alongside a transformer-based forecasting model.
+NVIDIA’s (Xu et al., 2024) research indicates that even when models already handle large context windows in text tasks, they can still achieve substantial performance gains by retrieving and using relevant data from external sources.
+To address the above challenges, we have identified a potential approach employing the concept of retrieval augmentation (RA).
+Concretely, we focus on implementing two central enhancements: a knowledge base schema that can efficiently index all historical series, and a cross-attention module embedded in a transformer model to integrate historical information for pinpointing and exploiting the most predictive segments, thereby refining prediction accuracy.
+The synergy between TSKB and RACA constitutes the RATSF framework, which significantly improves the performance of most univariate time series forecasting tasks.
+Currently, RATSF has been integrated into four key service areas within Fliggy, encompassing hotel bookings and after-sales services, train ticketing, and flight reservations and modifications.
+This integration has led to a notable decrease in forecasting errors, from an approximate range of 15% to approximately 8% across all the aforementioned sectors, which has significantly reduced personnel management costs, although the exact extent of cost savings cannot be divulged for commercial reasons.
+In concise terms, our main contributions are:
+1.
+We present a straightforward and manageable design for a time-series knowledge base (TSKB) based on the characteristic that time series data is easy to be stored in a structured way, which significantly facilitates efficient management of historical data.
+2.
+We introduce a versatile cross-attention module, retrieval augmented cross-attention (RACA) module, designed to integrate retrieved historical data into the forecasting process.
+This module is easily adaptable and can be seamlessly integrated with various time-series transformers.
+3.
+We raise a Retrieval-Augmented Time Series Forecasting (RATSF) framework, which decreases 7% forecasting errors in the real service areas within Fliggy and therefore reduced huge personnel practical management costs.
+Extensive experiments on more publicly available datasets demonstrate that our method achieves the best performance for the univariate time series forecasting task and is general for a broader range of industrial applications.
+2. Review
+Despite intense competition (Salinas et al., 2020; Das et al., 2023; Ye et al., 2022; Han and Woo, 2022), Transformer models (Vaswani et al., 2017) and their improved variants (Qi et al., 2021; Woo et al., 2022; Liu et al., 2021; Kitaev et al., 2020; Lim et al., 2021) have become the mainstream choice in time-series forecasting tasks.
+However, the computational complexity of the original Transformer model scales as O(n²) with respect to sequence length, significantly limiting the maximum sequence length it can handle and thereby constraining its ability to use historical data.
+Improvements in Time-Series Forecasting with Transformers.
+Numbers of works focus on enhancing the Transformer model’s capability to extract temporal features, thereby improving prediction accuracy, exemplified by Fedformer’s (Zhou et al., 2022) introduction of a frequency-augmented Attention mechanism that directly incorporates Fourier operators, a similar path taken by TimesNet (Wu et al., 2023).
+In contrast, Autoformer (Wu et al., 2021) proposes an operator that decomposes time series information into trend and periodic components.
+These two types of solutions perform well for relatively stationary sequences but may see performance drop when dealing with highly non-stationary time series data.
+Another core strategy involves learning a robust representation of the time series first, which is then used to enhance the forecasting accuracy.
+TNC (Tonekaboni et al., 2021) harnesses the core concept of contrastive learning and employs samples within a specific temporal neighborhood within a window as positive pairs, while treating samples from differing temporal neighborhoods as negative pairs.
+Despite these improvements having collectively boosted the ability of Transformers in time-series forecasting tasks, none has fundamentally increased the Transformer’s receptive field.
+Retrieve Augmented Method in NLP.
+In last two years, the Retrieval-Augmented Generation (RAG) approach (Guu et al., 2020; Borgeaud et al., 2022; Pasupat et al., 2021; Zhang et al., 2022; Shuster et al., 2021; Khandelwal et al., 2020; Zuo et al., 2019) has gained widespread adoption in the field of NLP.
+REALM harnesses a knowledge retriever to distill information from vast corpora and thereby enhance the performance of pre-trained language models.
+Transformer-XL+kNN (Bonetta et al., 2021) incorporates a K-Nearest Neighbors algorithm to search through training data, refining dialogue generation capabilities.
+In the context of named entity recognition tasks, U-RaNER (Tan et al., 2023) utilizes multi-modal heterogeneous retrieval techniques to boost knowledge retrive and, by integrating retrieved knowledge into the model, strengthens its understanding of queries and improves entity recognition accuracy.
+Retrieval Augmented Method in Time-series Forecasting.
+MQ-ReTCNN (Yang et al., 2022) is designed for complex time series prediction tasks involving multiple entities and variables.
+It employs a scoring function to compile relevant contexts from offline data, selects scored segments, and appends them to the prediction sequence.Its retrieval mechanism emphasizes leveraging historical sequences of one entity to enhance predictions about another, without directly utilizing historical data for direct prediction assistance.
+ReTime (Jing et al., 2022) creates a relation graph based on temporal closeness between sequences and employs relational retrieval instead of content-based retrieval.It does not optimally use historically similar sequences as reference points due to its inherent design limitations.
+Both MQ-ReTCNN and ReTime incorporate retrieval enhancement strategies but have yet to introduce a general and efficient retrieval technique specifically for single-variable time series prediction scenarios.
+Figure 1.
+RATSF uses a TSKB to store and index historical sequences, alongside a transformer-based forecasting model.
 As illustrated in the bottom-right corner, TSKB segments the full history for efficient retrieval.
@@
 RACA in the Decoder then merges $\mathbf{X}_{r}$ and $\mathbf{X}_{o}$ info to deliver result.
-In concise terms, our main contributions are:
-1. We offer a general retrieval-augmented time-series forecasting strategy, RATSF, that can be implemented across a variety of Transformer models and their derivatives to enhance performance.
-2. We employ a novel embedding strategy for precise historical piece retrieving.
-3. We present a straightforward and manageable design for a time-series knowledge base, which significantly facilitates efficient management of historical data.
-4.
-The method presented in this work currently serves as the operational algorithm for Fliggy’s service volume management system across all business domains, supporting not only hotel online bookings but also support businesses such as visa processing, flight and train ticket reservations, as well as entrance tickets to resort areas.
-Our method has demonstrated exceptional effectiveness in various practical application scenarios.
-2 Review
-Despite intense competition [20, 6, 30, 8], Transformer models [25] and their improved variants [19, 26, 16, 13, 14] have become the mainstream choice in time-series forecasting tasks.
-However, the computational complexity of the original Transformer model scales as O(n²) with respect to sequence length, significantly limiting the maximum sequence length it can handle and thereby constraining its ability to use historical data.
-Improvements in Time-Series Forecasting with Transformers.
-Numbers of works focus on enhancing the Transformer model’s capability to extract temporal features, thereby improving prediction accuracy, exemplified by Fedformer’s [33] introduction of a frequency-augmented Attention mechanism that directly incorporates Fourier operators, a similar path taken by TimesNet [27].
-In contrast, Autoformer [28] proposes an operator that decomposes time series information into trend and periodic components.
-These two types of solutions perform well for relatively stationary sequences but may see performance drop when dealing with highly non-stationary time series data.
-Another core strategy involves learning a robust representation of the time series first, which is then used to enhance the forecasting accuracy.
-TNC [24] harnesses the core concept of contrastive learning and employs samples within a specific temporal neighborhood within a window as positive pairs, while treating samples from differing temporal neighborhoods as negative pairs.
-Despite these improvements having collectively boosted the ability of Transformers in time-series forecasting tasks, none has fundamentally increased the Transformer’s receptive field.
-Retrieve Augmented Method in NLP.
-In last two years, the Retrieval-Augmented Generation (RAG) approach [7, 4, 18, 31, 21, 12, 34] has gained widespread adoption in the field of NLP.
-REALM harnesses a knowledge retriever to distill information from vast corpora and thereby enhance the performance of pre-trained language models.
-Transformer-XL+kNN [3] incorporates a K-Nearest Neighbors algorithm to search through training data, refining dialogue generation capabilities.
-In the context of named entity recognition tasks, U-RaNER [22] utilizes multi-modal heterogeneous retrieval techniques to boost knowledge retrive and, by integrating retrieved knowledge into the model, strengthens its understanding of queries and improves entity recognition accuracy.
-Retrieval Augmented Method in Time-series Forecasting.
-MQ-ReTCNN [29] is designed for complex time series prediction tasks involving multiple entities and variables.
-It employs a scoring function to compile relevant contexts from offline data, selects scored segments, and appends them to the prediction sequence.Its retrieval mechanism emphasizes leveraging historical sequences of one entity to enhance predictions about another, without directly utilizing historical data for direct prediction assistance.
-ReTime [11] creates a relation graph based on temporal closeness between sequences and employs relational retrieval instead of content-based retrieval.It does not optimally use historically similar sequences as reference points due to its inherent design limitations.
-Both MQ-ReTCNN and ReTime incorporate retrieval enhancement strategies but have yet to introduce a general and efficient retrieval technique specifically for single-variable time series prediction scenarios.
-Figure 2: Each RACA has two cross-attention modules: one utilizes the Encoder output as K,V, while the other employs embedded retrieved sequences as K,V.
-Both outputs are concatenated and passed through a Linear module to reshape back to the input dimensions.
-3 Method
-Subsequently, we first provide an overview of the RATSF’s overall structure and then examine the key features of each component in detail.
-The simplest description of RATSF is to first build a TSKB and then retrieve relevant sequences from it to enhance the forecasting model.
-For a Time Series Forecasting (TSF) task within a specific context, we first construct a TSKB using the schema detailed in Section 3.2.
-Next, we utilize a retrieval method to gather relevant sequences, with this method being introduced in Section 3.3.
-Finally, we integrate these retrieved sequences into the forecasting model, with its core computational component, RACA, being introduced in Section 3.1.
-Figure 1 shows the data flow of RATSF.
-3.1 RACA
-As previously mentioned, Retrieval-Augmented Cross-Attention (RACA) is a module designed to be plugged-into a time-series forecasting model.
-In our demonstration, we employ the vanilla Transformer as the time-series forecasting model; however, it’s important to note that the RACA module is designed to be compatible and can be seamlessly plug into any transformer-based time-series model.
-we now provide an explanation of the abbreviations represented by subscript and superscript symbols in the following paragraphs: The letter ’o’ for original, represents the input time-series piece entering the encoder; letter ’b’ for embedding , letter ’e’ for encoder; letter ’r’ for retrieved sequence, letter ’f’ for sequence for forecasting; ’l’ for layer ; ’D’ for hidden dimension.
-In transformer architecture, an encoder consists of an embedding module and multiple self-attention modules.
-Initially, the raw time series data $\mathbf{X}_{o}=\left\{x_{1},...,x_{t}\right\}$, $\mathbf{X}_{o}$ when processed through the embedding module, generates an output as the $\mathbf{X}^{b}_{o}$; subsequently, the output resulting from all the encoder calculation is referred as the $\mathbf{X}^{e}_{o}$.
-RACA is introduced within the decoder part: besides conventionally utilizing the encoder output and the place holding embedding for the forecasting sequence, it incorporates information from N most relevant history sequences retrieved from TSKB.
-Like shown in Figure 2, these retrieved sequences are transformed through the encoder’s embedding module, generating Ns $\mathbf{X}^{e}_{r}$  [^1] In section 4, when comparing four Transformer models horizontally, the impact of the Encoder extends to both representation retrieval and decoding stages.
-To isolate the contribution of the Encoder to the retrieval performance, we employ the Encoder output as the retrieval scheme while incorporating the representation $\mathbf{X}^{b}_{r}$ rather than $\mathbf{X}^{e}_{r}$ into RACA.
-In section 3.3, we directly compare the performance of using DTW alone versus Encoder output for retrieval.
-When integrated into the Decoder, the Encoder output’s affecting both retrieval effectiveness and decoding quality.
-Given the significant differences between Encoders trained solely with DTW and those following our proposed method,as shown in B.2.
-To ensure consistency across experiments, $\mathbf{X}^{e}_{r}$ is consistently replaced with $\mathbf{X}^{b}_{r}$ in this paper’s all experiments., and we concatenate these N vectors into one with shape$[1,N*L_{r},D]$ for later use.
-In Transformer-based time series forecasting models, placeholders are embedded into a $[1,L_{f},D]$ matrix, $\mathbf{H}_{f}$.
-The decoder consists of $l$ stacked RACA modules for inputs of length $L_{f}$.
-Each module has two parallel units: Unit 1, like in Function (1), uses $\mathbf{H}_{f}$ as Query with Key and Value from $\mathbf{X}^{e}_{o}$, outputting $\mathbf{H}^{l,1}_{f}$ of the same shape.
-Unit 2, retrieval-augment part, like in Function (2) ,also queries $\mathbf{H}_{f}$ but fetches Keys and Values from the concatenated sequence of $\mathbf{X}^{b}_{r}$, generating $\mathbf{H}^{l,2}_{f}$.
-$s$ in the Functions is the scaling factor.
-$\mathbf{H}^{l,1}_{f}=\text{softmax}\left(\frac{\mathbf{H}_{f}\mathbf{X}^{e}_{o}}{s}\right)\mathbf{X}^{e}_{o},$ (1)
-$\mathbf{H}^{l,2}_{f}=\text{softmax}\left(\frac{\mathbf{H}_{f}\mathbf{X}^{e}_{r}}{s}\right)\mathbf{X}^{e}_{r},$ (2)
-$\mathbf{H}^{l+1}_{f}=\delta\left(\text{concat}\left(\mathbf{H}^{l,1}_{f},\mathbf{H}^{l,2}_{f}\right)\right).$ (3)
-As shown in Function 3, after concatenating both intermediate vectors,$\mathbf{H}^{l,1}_{f}$ and $\mathbf{H}^{l,2}_{f}$, along the sequence dimension, forms a $[1,2*L_{f},D]$ vector, it undergoes linear transformation before being passed to the next layer.
-3.2 TSKB
-In the development of TSKB, we employ an innovative strategy for constructing the indexing sequence (K) and content sequence (V).
-Unlike traditional methods that solely store and retrieve the entire original sequence, TSKB preserves the original sequence as the core content while selects certain segments from it to construct a discriminative indexing sequence.
+3. Method
+3.1. Setting & Notations
+Before elucidating our approach, we first clarify the setting of our problem and define some symbols that will be used throughout the text.
+In actual business operations, service volume is influenced by a multitude of factors, some of which are challenging to fully represent through variables.
+To simplify the reasoning process and enhance the versatility of our system, we have configured our hotel service volume prediction using a uni-variant time-series forecast setup.
+This means that the inputs to the model consist solely of the the values of the time series and it’s temporal features .
+Due to the requirements of our actual task—where personnel management necessitates advance preparation for recruitment, short-term scheduling, and training.
+Our forecasting task is designed to start from a specific time point $t$ and predict a sequence of $L_{f}$ data points in one go.
+Time Marking.
+We take moment $t$ as the reference point, and the collection of the future time points to be forecasted is represented as ${\left[t+1,t+2,\ldots,t+l_{f}\right]}$, where $l_{f}$ denotes the length of the forecast period.
+Concurrently, we define the retrieve segment length as $l_{r}$, with the indexing sequence $\mathbf{K}$ in the TSKB having a length of $L_{r}$, and the length of the V sequence as $l_{v}$.
+Sequence Source Marking.
+The subscript o denotes the original sequence, that is, the sequence $\mathbf{X_{o}}={x_{1},...,x_{t}}$ prior to moment $t$; $\mathbf{X_{f}}$ represents the sequence to be forecasted; $\mathbf{X_{r}}$ signifies the $\mathbf{V}$ sequence retrieved from the TSKB.
+Transformer Processing Marking.
+The superscript is used to indicate the results after processing by various parts of the Transformer.
+$\mathbf{X^{b}_{o}}$ denotes the original sequence after Embedding, and $\mathbf{X^{e}_{o}}$ denotes the sequence after the Encoder.
+Within the Decoder, H represents the internal hidden state of the Decoder, $l$ represents the layer number, and $H^{l}$ indicates the hidden state processed by the $l$-th layer.
+3.2. Overview of RATSF
+The RATSF system is composed of two core components: a Time Series Knowledge Base (TSKB, detailed in Section 3.3) and a Transformer-based time series forecasting model.
+The latter has been enhanced by replacing its original decoder with our novel Retrieval Augmented Cross-Attention (RACA, described in Section 3.4) module.
+The TSKB efficiently segments and archives historical time series data, establishing a precise indexing system.
+The transformer model, with the integration of RACA, effectively utilizes retrieved historical data to significantly enhance the accuracy of its predictions.
+The data flow diagram of RATSF is illustrated in Figure 1, providing a visual representation of the process.
+3.3. TSKB
+Unlike traditional methods that solely store and retrieve the entire original sequence, TSKB preserves the original sequence as the core content (V) while selects certain segments from it to construct a discriminative indexing sequence (K).
 This approach can be likened to processing a written piece where conventional methods involve direct full-text searches for desired information, which may be inefficient and less precise; whereas with TSKB, it’s akin to extracting key headlines from the body of the text to serve as indices that facilitate rapid access to core information.
 This dual-sequence structure allows the system, when handling large-scale data, to efficiently locate and access targeted portions of the original sequence through customized index sequences, thereby significantly enhancing overall performance.
-In this article, to illustrate our system design, we employ a fixed slicing logic to select segments from V for computing the retrieval representation.
-In section 3.3, we will present in detail how the representation of V is learned.
-Figure 3: Within the cold start stage, we use raw sequence of retrieval input to find Top-K relevant sequences using DTW.
+3.3.1. Sequential Slicing
+As illustrated in Figure 2, the content sequence V is obtained using a rolling window approach with a step size of $S$ and a window length of $L_{v}$.
+This approach allows us to sample all historical sequences and incorporate them into TSKB.
+Since prediction models only use present data to infer the future, we align indexing sequence K with this constraint.
+We extract the initial segment of each V, having a length of $L_{r}$, to serve as its index K.
+The selection of $L_{v}$, $L_{r}$ and $S$ is introduced in section 3.5.
+Figure 2. TSKB utilizes a rolling window of length $L_{v}$ to collect V, with an indexing segment of length $L_{r}$ taken from its leading part as K , and advances the window in steps of size S.
+Figure 3.
+Within the cold start stage, we use raw sequence of retrieval input to find Top-K relevant sequences using DTW.
 After one epoch, we switch to Euclidean Distance for retrieval embedding to gather Top-K matches.
-Figure 4: TSKB utilizes a rolling window of length $L_{v}$ to collect V, with an indexing segment of length $L_{r}$ taken from its leading part as K , and advances the window in steps of size S.
-3.3 Learning Retrieval Embedding
-Judging the quality of a retrieval representation mainly depends on how well it captures similarities at key information points relevant to forecasting tasks.
+3.3.2. Embedding Learning
+A key approach to enhancing the recall and accuracy of knowledge base retrieval is to employ a well-trained embedding vector to improve the precision of the index.
+In the following sections, we will focus on how to train and utilize these index embedding vectors.
+Judging the quality of a retrieval embedding mainly depends on how well it captures similarities at key information points relevant to forecasting tasks.
 Time series data is complex due to numerous information points and the challenge of presetting comparative weights.
 Therefore, we use this principle: if a representation closely mirrors the important details needed for prediction, its overall performance will be better.
-Later, we’ll explain the exact steps to put this principle into practice.
-We acquire the retrieval embedding for TSKB through two procedures.
-3.3.1 Sequential Slicing
-To obtain content sequence V, we utilize a rolling window approach where the step size is $S$ and the window length is $L_{v}$, to sample all historical sequences and incorporate them into TSKB.
-Since prediction models only use present data to infer the future, we align indexing sequence K with this constraint.
-We extract the initial segment of each V, having a length of $L_{r}$, to serve as its K.
-The selection of $L_{v}$, $L_{r}$ and $S$ is introduced in section 4.
-3.3.2 Embedding Learning
-To guarantee the embedding in capturing the essential information for forecasting, we utilize the encoder of the RATSF’s forecasting model.
-The model inherently excels at selecting precise information during training to improve forecast accuracy.
-In practice, every indexing sequence K from the knowledge base undergoes the RATSF forecasting model’s encoder layer to generate a retrieve embedding.
+To ensure that the embeddings capture the essential information for forecasting, we employ the encoder from the RATSF’s forecasting model, which is designed to select precise information during training, thereby enhancing forecast accuracy.
+Specifically, each indexing sequence K rom the knowledge base is processed through the encoder layer of the RATSF forecasting model to generate a retrieval embedding.
 Meanwhile, in the early stages of training the forecasting model, its encoder is not yet capable of producing representations that accurately match target sequences.
-The effectiveness and learning pace of the RATSF’s forecasting model are significantly correlated with whether or not it retrieves historical sequences that are beneficial to the forecasting task.
-To avoid time and data-consuming iterations stemming from a random initialization state, we introduce DTW as an auxiliary tool during the initial phase of model training, as shown in left part of Figure 3.
-Initially, DTW is utilized for similarity-based sequence retrieval, which is then incorporated into the iterative training process of the RATSF forecasting model.
-After finish one epoch of training, we transition to using embedding generated by the forecasting model itself for similar sequence retrieval, continuing the training until the model converges.
-Figure 5: The lower plot displays the top 3 relevant retrieved sequences for one sample that are end-to-end connected.
-Each retrieved sequence has a length of 21 ($L_{f}$), comprising 14 points preceding the forecasting date and 7 points to be forecasted, highlighted with an orange background.
-The color bars represent aggregated cross-attention values at the first forecasting point, in which lighter color indicate higher attention value.
-In the upper part of the figure, predicted (yellow) and ground truth (orange) values align at corresponding positions with retrieved sequence.
-For this sample, the model notably focuses on segments indicating ascent and predicts a subsequent increase in value.
-Table 1: Performance comparison of four Transformer models with and without the application of RATSF on three experimental datasets.
-All results are evaluated using MSE (Mean Squared Error) and MAE (Mean Absolute Error), and the best result is bolded.
-models transformer nstransformer autoformer fedformer
-metric mse mae mse mae mse mae mse mae
-FHSV baseline 4344450.512 1349.146 4414326.000 1242.128 6096673.500 1752.798 5517485.000 1638.275
-RATSF 3421546.027 1101.172 3823635.000 1070.933 5597297.500 1657.561 5386818.500 1595.657
-ETTh1 baseline 3.187 1.391 1.501 0.920 8.651 2.264 4.159 1.597
-RATSF 2.366 1.180 1.407 0.886 7.300 2.113 3.952 1.565
-Exchange baseline 0.00165 0.03134 0.00014 0.00922 0.00058 0.01906 0.00172 0.03493
-RATSF 0.00068 0.02012 0.00013 0.00912 0.00067 0.02000 0.00068 0.02084
-4 Experiments
-In order to evaluate the generalizability and efficacy of RATSF in industry applications, we initially carry out a range of experiments to contrast the performance changes among typical time-series models when augmented with RATSF across various tasks, focusing particularly on the Fliggy Hotel Service Volume Dataset (FHSV).
-Following this, we employ a sequence of ablation studies to separately demonstrate the design choice of each improvement – RACA, TSKB, and Retrieval Embedding.
-4.1 Experiment Setup
+The model’s effectiveness and learning pace are significantly correlated with its ability to retrieve historical sequences beneficial to forecasting tasks.
+To avoid time and data-consuming iterations stemming from a random initialization state, we introduce Dynamic Time Warping (DTW(Müller, 2007)) as an auxiliary tool during the initial phase of model training, as shown in left part of Figure 3.
+DTW is initially used for similarity-based sequence retrieval, aiding in the iterative training process of the RATSF forecasting model.
+After completing one epoch of training, we transition to using embeddings generated by the forecasting model itself for sequence retrieval, continuing the training until the model converges.
+Figure 4.
+Each RACA has two cross-attention modules: one utilizes the Encoder output as K,V, while the other employs embedded retrieved sequences as K,V.
+Both outputs are concatenated and passed through a Linear module to reshape back to the input dimensions.
+3.4. RACA
+As previously mentioned, Retrieval-Augmented Cross-Attention (RACA) is a module designed to be integrated into a time-series forecasting model, coupled with the TSKB.
+In our demonstration, we employ the vanilla Transformer as the time-series forecasting model; however, it’s important to note that the RACA module is designed to be compatible and can be seamlessly plug into any transformer-based time-series model.
+We will specifically demonstrate this in Section 4 .
+The decoder consists of $l$ stacked RACA modules for inputs of length $L_{f}$.
+Each module has two parallel units: Unit 1, like in Function (1), uses $\mathbf{X}^{b}_{f}$ as Query with Key and Value from $\mathbf{X}^{e}_{o}$, outputting $\mathbf{H}^{l,1}_{f}$ of the same shape.
+Unit 2, retrieval-augment part, like in Function (2) ,also queries $\mathbf{X}^{b}_{f}$ but fetches Keys and Values from the concatenated sequence of $\mathbf{X}^{e}_{r}$, generating $\mathbf{H}^{l,2}_{f}$.
+$s$ in the Functions is the scaling factor.
+Like shown in Figure 4, retrieved sequences are transformed through the encoder’s embedding module, generating Ns $\mathbf{X}^{e}_{r}$, and we concatenate these N vectors into one with shape$[1,N*L_{v},D]$ for later use.
+(1) $\mathbf{H}^{l,1}_{f}=\text{softmax}\left(\frac{\mathbf{X}^{b}_{f}\mathbf{X}^{e}_{o}}{s}\right)\mathbf{X}^{e}_{o},$
+(2) $\mathbf{H}^{l,2}_{f}=\text{softmax}\left(\frac{\mathbf{X}^{b}_{f}\mathbf{X}^{e}_{r}}{s}\right)\mathbf{X}^{e}_{r},$
+(3) $\mathbf{H}^{l+1}_{f}=\delta\left(\text{concat}\left(\mathbf{H}^{l,1}_{f},\mathbf{H}^{l,2}_{f}\right)\right).$
+As shown in Function 3, after concatenating both intermediate vectors,$\mathbf{H}^{l,1}_{f}$ and $\mathbf{H}^{l,2}_{f}$, along the sequence dimension, forms a $[1,2*L_{f},D]$ vector, it undergoes linear transformation before being passed to the next layer.
+3.5. Deployments Procedure
+In this section, we briefly introduce the procedure to deploy RATSF in a new field.
+3.5.1. TSKB Initialization
+Set the initial retrieval length ($L_{r}$), which can be equivalent to the prediction length ($L_{f}$).
+Set the sequence length for the V sequence ($L_{v}$) to be at least the sum of $L_{r}$ and $L_{f}$.
+This ensures adequate info for retrieval and forecast.
+It is suggested to set several $L_{v}$ values, such as $L_{r}$ + $L_{f}$, 2*$L_{f}$ + $L_{r}$, 3*$L_{f}$ + $L_{r}$, etc.,to build several TSKBs for optimal value selection later.
+Set the rolling window step (stride) to 1 and collect $V$ sequences for each TSKBs.
+Then, Select the initial $l_{r}$ elements of $V$ sequences to form $K$ sequences.
+3.5.2. Identifying Optimal $L_{v}$
+Train diverse RATSF models on their respective TSKBs and evaluate each by prediction accuracy.Select the model with the highest prediction accuracy as the optimal RATSF model ($M^{*}$), and identify the corresponding V sequence length ($L_{v}$) as the final chosen value ($L_{v}^{*}$).
+3.5.3. Adjusting Retrieval Length $L_{r}$
+Adjust the retrieval length ($L_{r}$) to observe the impact on the performance of the optimal model (M*).
+The adjustment range could be from $0.5L_{f}$ to $L_{v}$, with each adjustment increment being $0.5L_{f}$.
+Record the $L_{r}$ value that enables M to achieve the highest prediction accuracy, denoted as $L_{r}^{*}$.
+3.5.4. TSKB and Model Optimization
+With the confirmed optimal retrieval length ($L_{r}*$) and V sequence length ($L_{v}^{*}$), reconfigure the TSKB. Reinitialize and train the RATSF model with the updated TSKB till converge.
+4. Experiments
+To demonstrate the superiority of our method in service volume forecasting, we first conducted experiments using the Fliggy Hotel Service Volume Dataset (FHSV), showcasing the performance enhancement of typical time-series models when integrated with RATSF.
+Additionally, to validate the general applicability of our approach, we extended our experiments to three other datasets.
+Furthermore, through a series of ablation studies, we individually assessed the impact of each key design choice in RATSF to confirm the correctness of our detailed selections.
+4.1. Experiment Setup
 Datasets.
 Given that our primary focus lies in the forecasting task within the context of customer service volume management, we initially employed the Fliggy Hotel Service Volume Dataset to substantiate the efficacy of RATSF, the detail description of the dataset is in Appendix A.
-Additionally, to demonstrate the performance of our approach in other contexts, we employed two classic datasets: ETT and Exchange.
-Models.We contrasted the vanilla Transformer against three of its prominent time-series domain variants: Fedformer, Autoformer, and NS-Transformer.
-Specifically, in keeping with the intrinsic sequential decomposition characteristics of Fedformer and Autoformer, historical segments were initially broken down using each model’s native calculations, for example the seasonal init and trend-cyclical init in Autoformer.
-Subsequently, these decomposed segments were embedded and fed into the RACA module for further processing.
-Forecasting Setting.Consistent with the original intent of validating the effectiveness of RATSF in the context of customer service volume management, we adopt the business setting from Fliggy.
+Additionally, to demonstrate the performance of our approach in other contexts, we employed three classic time-series forecasting datasets: ETT(Zhou et al., 2021), Exchange(Zhao et al., 2019) , Traffic(tra, 2017) .
+Models.In Section 3.4, we established that the Retrieval Augmented Cross-Attention (RACA) module is compatible with various Transformer variants.
+To underscore this versatility, we selected three prominent time-series domain models: Fedformer(Zhou et al., 2022), Autoformer(Wu et al., 2021), and NS-Transformer(Liu et al., 2022), and compared them to the vanilla Transformer.
+Leveraging the specialized operators within Fedformer and Autoformer that cater to time-series periodicity and trends, such as Autoformer’s seasonal and trend-cyclical initializations, we have seamlessly integrated these models with the RACA module.
+This integration allows for the effective incorporation of their distinct sequential decomposition capabilities, enhancing the overall processing power.
+Furthermore, to demonstrate the advantage of our recall mechanism in the RATSF framework, we compared it with ReTime, a RAG-based model that also uses specialized retrieval techniques.
+Since ReTime has not released their code and their application domain is distinct from our focus on univariate time-series forecasting, we reconstructed the Relational Retrieval method based on their published paper.
+(It is worth noting that, as of the time of writing this paper, neither MQ-ReTCNN nor ReTime have released their code to the public.)
+As shown in Figure 1, the training outcomes of the encoder’s output significantly influence both the retrieval quality and the decoder’s performance.
+In the main experiments, we will compare the retrieval schemes of RATSF and ReTime.
+Furthermore, in several subsequent experiments, we will elucidate the differences between RATSF’s retrieval approach and DTW.
+Consequently, for the majority of the experiments in this paper, to facilitate these comparisons, we have replaced the input to RACA from $\mathbf{X}^{e}_{r}$ to $\mathbf{X}^{b}_{r}$, and we will provide an ablation study to explore the specific effects arising from this substitution.
+Forecasting Setting.
+Like we claim in Section 3.1, we adopt the uni-variant time-series forecasting setting.
 In the Fliggy’s practical operations, we are required to finalize our forecasting for the upcoming week’s day-by-day staffing needs one week in advance, which is why the actual prediction window period is set at 7 days, represented as 7 tokens.
@@
 Regarding normalization, we employ unified $\mu$ and $\sigma$ values for whitening operations during the preprocessing stage and inverse normalization is carried out after forecasting.
-We have applied the same data processing scheme to both the ETT and Exchange datasets.
+We have applied the same data processing scheme to both the ETT, Exchange and Traffic datasets.
 Training Setting.We have selected the Adam optimizer, with the batch size$=$64 and max training epochs$=$10.
@@
 This configuration has been applied uniformly across all experiments.
-4.2 Main Restult
-Table 1 presents the performance comparison of four Transformer models with and without the application of RATSF on three experimental datasets.
-In the scenario of forecasting with the Fliggy Hotel Customer Service Volume Dataset, the Transformer model’s MAE loss dropped by 18% after incorporating RATSF, potentially translating to a reduction of approximately 200 redundant personnel cost in practical management operations.
-Notably, even with the NSTransformer [17], Autoformer, and FedFormer models, which have undergone temporal optimization in their attention operators, adopted the RATSF design, they respectively recorded decreases in MAE loss by 14%, 5%, and 4%.
+Table 1.
+Performance comparison of the Transformer models with their retrieval-based variants on four experimental datasets.
+’baseline’ denotes the original model without retrieval, compared to Relational Retrieva from ReTime and our approach RATSF.All results are evaluated using MSE (Mean Squared Error) and MAE (Mean Absolute Error), and the best result is bolded.
+models transformer nstransformer autoformer fedformer
+metric mse mae mse mae mse mae mse mae
+FHSV baseline 4344450.512 1349.146 4414326.000 1242.128 6096673.500 1752.798 5517485.000 1638.275
+Relational Retrieval 3737283.320 1276.836 4037294.000 1168.127 5821383.300 1703.328 5411003.700 1610.320
+RATSF 3421546.027 1101.172 3823635.000 1070.933 5597297.500 1657.561 5386818.500 1595.657
+ETTh1 baseline 3.187 1.391 1.501 0.920 8.651 2.264 4.159 1.597
+Relational Retrieval 2.876 1.206 1.465 0.900 8.203 2.197 3.012 1.572
+RATSF 2.366 1.180 1.407 0.886 7.300 2.113 3.952 1.565
+Exchange baseline 0.00165 0.03134 0.00014 0.00922 0.00058 0.01906 0.00172 0.03493
+Relational Retrieval 0.00143 0.02845 0.00013 0.00920 0.00069 0.02013 0.00102 0.02821
+RATSF 0.00068 0.02012 0.00013 0.00912 0.00067 0.02000 0.00068 0.02084
+Traffic baseline 0.00007 0.00564 0.00006 0.00569 0.00008 0.00643 0.00012 0.00876
+Relational Retrieval 0.00007 0.00534 0.00006 0.00512 0.00007 0.00601 0.00009 0.00718
+RATSF 0.00006 0.00498 0.00006 0.00472 0.00007 0.00514 0.00007 0.00650
+4.2. Main Restult
+Table 1 displays the prediction accuracy of the Transformer and its variants on four datasets, comparing performance without RA, with ReTime’s RA, and with our RATSF approach.
+In the Fliggy Hotel Customer Service Volume Dataset forecasting scenario, the Transformer model’s MAE loss was reduced by 5.34% with ReTime’s Relational Retrieval and by 18% after incorporating RATSF.
+This improvement could potentially translate to a reduction of approximately 200 in redundant personnel costs in practical management operations.
+A horizontal analysis of the first row in Table 1 reveals that all compared Transformer-based time-series model variants experienced a significant reduction in MAE after employing Retrieval-Augmented (RA) strategies.
+Specifically, the adoption of RATSF resulted in respective decreases of 14%, 5%, and 4%.
+In contrast, the application of ReTime’s retreval method led to improvements of 5.9%, 2.8%, and 1.7% compared to their original designs without RA.
+These results clearly demonstrate two conclusions: first, effective RA strategies can enhance the performance of time-series Transformer models; second, RATSF exhibits a significant advantage in its retrieval strategy.
 Furthermore, cross-comparison reveals that Autoformer and FedFormer yield slightly inferior results compared to the Transformer, mainly due to the stronger non-stationarity inherent in customer service volumes data within the hotel industry.
@@
 In such instances, historical data from periods close to other short holidays provide more valuable insights than simple temporal periodicity and short-term trends; specific case studies will be showcased later.
-Additionally, across the ETT and Exchange datasets, the adoption of RATSF design in the Transformer and its time-series optimized variants led to noticeable reductions in MSE metrics.
+Additionally, across the ETT, Exchange and the Traffic dataset, the adoption of RATSF design in the Transformer and its time-series optimized variants led to noticeable reductions in MAE metrics.
 Moreover, among these, the NSTransformer consistently outperforms the other models, likely because its mean-adapted feature proves effective in a broad range of time-series prediction domains.
-4.3 Ablation Study
-Section 4.3.1 comparatively analyzes the differences in various designs for integrating historical pieces into forecasting.
-Section 4.3.2 examines the impact of integrating varying numbers of historical pieces on forecast effectiveness.
+4.3. Ablation Study
+We confirm the effectiveness of several design choices in this section.
+Section 4.3.1 emphasizes the superiority of RACA’s approach in integrating historical data for improved forecasting.
 Section 4.3.2 demonstrates the advantages of utilizing model encoder for recalling historical pieces.
-Table 2: Ablation of the specifically designed structures within RACA on FHSV scenario.
+Section 4.3.3 demonstrates the benefits of incorporating DTW into the training process.
+Figure 5.
+This figure shows RACA’s use of retrieved sequences for a random forecast sequence $\mathbf{X_{f}^{*}}$.The lower half features the top 3 similar sequences of $\mathbf{X_{f}^{*}}$, concatenated along the time axis to form RACA’s input.
+Attention weights at the first forecast point ($x_{t+1}$) are marked by data bars, with yellow and green indicating RACA’s focus on downward-curving segments.
+The upper half replicates prediction results (yellow ”pred” lines for $x_{t+1}...x_{t+f}$) and true values (orange ”gt” lines), aligned with the corresponding final $L_{f}$ lengths of the retrieved sequences.
+This visual comparison highlights the model’s pattern recognition, with RACA’s focus areas correlating to the predictions’ inflection at $x_{t+1}$.For this sample, the model notably focuses on segments indicating ascent and predicts a subsequent increase in value.
+Table 2. Ablation of the specifically designed structures within RACA on FHSV scenario.
 dataset methods metric
@@
 baseline mae 1349.146
-baseline + mse 4071384.75
-enhanced encoder mae 1226.850
+baseline + mse 4071384.750
+Design One mae 1226.850
 mse 3421546.027
 FHSV baseline + RACA mae 1101.172
-4.3.1 RACA’s design choice
+4.3.1. Integrate Historical Sequences with RACA
 Based on the Transformer architecture, two designs are considered for integrating historical pieces into the forecasting model: Design One represents a straightforward approach where historical pieces are combined with the forecasting context inputs within the encoder; Design Two employs our advanced RACA design that integrate historical pieces using cross-attention mechanisms within the decoder.
@@
 A deeper examination disclosed that the Cross Attention mechanism in the RACA design efficiently guides the forecast sequence to concentrate on past trends analogous to the current scenario.
-Figure 5 illustrates the attention map within the RACA for the data points under forecast, where highlighted color blocks indicate that the data point being forecast places significant attention on historical segments that exhibiting similar trend patterns, and these historical segments’ trends closely align with the actual waveform following the data point in the ground truth.
-4.3.2 Optimal Retrieval Count
-Table 3: Results of recalling different numbers of pieces in our RATSF model.
-dataset methods metric
-mse 4344450.512
-FHSV baseline mae 1349.146
-mse 3678230.020
-recalled Top-1 pieces mae 1160.101
-mse 3700934.754
-recalled Top-2 pieces mae 1155.952
-mse 3421546.027
-recalled Top-3 pieces mae 1101.172
-mse 3460814.538
-recalled Top-4 pieces mae 1108.895
-mse 3750426.751
-recalled Top-5 pieces mae 1231.564
-Having established that integrating effective historical pieces positively impacts forecasting, we then inquire about the optimal number of sequences to retrieve.
-To demonstrate the effect of this choice, we gradually increase the number of integrated retrieve sequences from 0 to 5.
-As shown in Table 3, during the process of expanding from integrating no historical piece to integrating up to three historical pieces, the MAE error consistently decreases; however, when the quantity of integrated historical pieces is further increased, the forecasting error starts to escalate.
-Upon analyzing samples, the insight is obvious: not all of the top five retrieved historical piece for most of samples closely resemble the Ground Truth.
-The reason being, focusing solely on the TopK ranking without adequately considering similarity thresholds can easily lead to a scenario where, as K increases, the quality of the retrieved pieces becomes less assured and more prone to introduce noise and irrelevant information, thus potentially causing confusion instead.
-4.3.3 Retrieval Embedding
-We claimed that DTW is not optimal as the ultimate retrieving method.
-In this part of ablation we show that with an experiment on FHSV.
-We compare DTW, a trained MLP and current design choice, the forecasting model’s encoder in terms of accuracy.
-Table 4 clearly demonstrates that the forecasting model’s encoder output provide embedding outperforms both the other two schemes significantly.
-The advantages of encoder output over MLP are evident, particularly because the attention mechanism within the transformer’s encoder is better suited to discern and refine the interdependence’s among data points in sequences, thereby abstracting more accurate patterns.
-Table 4: Comparison of several variants of retrieval representation schemes.
-dataset methods metric
+In Figure 5, we illustrate how retrieved sequences influence the predicted values for the first forecast point of a specific sequence through RACA’s cross-attention mechanism.
+The lower half features the top 3 similar sequences, concatenated along the time axis to form RACA’s input.
+Attention weights at the first forecast point ($x_{t+1}$) are marked by data bars, with yellow and green indicating RACA’s focus on downward-curving segments.
+The upper half replicates prediction results (yellow ”pred” lines for $x_{t+1}...x_{t+f}$) and true values (orange ”gt” lines), aligned with the corresponding final $L_{f}$ lengths of the retrieved sequences.
+This visual comparison highlights the model’s pattern recognition, with RACA’s focus areas correlating to the predictions’ inflection at $x_{t+1}$.
+Table 3. Comparison of several variants of retrieval representation schemes on the FHSH dataset.
+stage methods metric
 mse 4344450.512
@@
 mse 3421546.027
-FHSV baseline + encoder mae 1101.172
-5 Conclusion
-In this paper, we propose the RATSF and explain its effectiveness in enhancing time-series forecasting performance.
-Specifically, our approach improves from three dimensions: first, we integrate the RACA mechanism into the transformer model architecture, which leverages historical data segments similar to current contexts to bolster forecasting prowess.
-Second, we construct a KV schema for a time-series knowledge base that significantly enhances the flexibility and precision of the retrieval process.
-And third, we utilize the Encoder component within the forecasting model to embed K sequences, thereby refining the accuracy of retrieving relevant historical information.
-The method introduced in this study has been actively operational within Fliggy’s service volume management system for several months.
-Of particular importance is that due to the strong flexibility and scalability inherent in this design, the method can play a significant role across a wide array of practical application scenarios.
-Appendix A Fliggy Hotel Service Volume Dataset
-The Fliggy Hotel Service Volume Dataset (FHSV) is derived from service records within the Fliggy APP’s customer service center, capturing the volume of daily customer service operations related to inquiries about hotel reservations, cancellations, and changes.
-Each instance of a service request received either through the APP interface or by phone is counted as a single service event.
-Daily aggregates of such events are recorded, forming individual data points.
-This dataset comprises 1740 data points collected between January 1, 2019, and October 7, 2023.
-To facilitate experimentation, this time series dataset has been sequentially partitioned in a chronological order into three subsets: the train set (from January 1, 2019, up until May 24, 2022), the evaluation set (from May 25, 2022, to October 31, 2022), and the test set (from November 1, 2022, onward).
-The reported experimental results are based on the model’s performance on the test set alone.
-Appendix B Other Experiment Details
-B.1 Length of Retrieval Index Segment of Knowledge Base
-Table 5: Length of Retrieval Index Segment of Knowledge Base.
-dataset retrieval length metric
-mse 4621934.067
-4 mae 1303.329
-mse 3828479.700
-7 mae 1178.574
+representation retrieval baseline + encoder mae 1101.172
 mse 3421546.027
-14 mae 1101.172
-mse 3747974.421
-21 mae 1208.379
-mse 3816283.253
-28 mae 1243.133
-mse 3913793.800
-FHSV 35 mae 1332.472
-The length of the retrieval segment in a knowledge base significantly affects the accuracy of retrieving the original sequence.
-Through controlled experiments, we identified the most suitable retrieval segment length for the Fliggy scenario.
-As shown in Table 5, we incrementally extended the length of the retrieval segment K from 4 units to 35 (corresponding to half to five times the length of the prediction sequence at 7 units), then observed the impact on model performance as measured by MSE and MAE.
-The results revealed that during the expansion of the retrieval segment from 4 to 14 units, the predictive performance improved.
-However, beyond this point, as the retrieval segment continued to lengthen, the predictive performance began to decline.
-This trend demonstrates that, in practical scenarios, as the retrieval segment grows from short to long, its representation transitions from being information-poor to increasingly mixed and complex.
-This situation underscores why, in traditional time series databases, using full-ordered representations for retrieval often yields effective outcomes.
-By contrast, our K,V design effectively circumvents this issue, providing flexible and accurate retrieval results.
-B.2 Effectiveness of Using DTW as Auxiliary Training
-In the article, we claimed the effectiveness of using DTW as a feature retrieval technique during the cold start phase to aid in single-epoch training for forecasting models.
-To validate the necessity of this approach, we designed a set of comparative experiments: a baseline group without DTW-assisted training (named by end to end), and other groups that were each trained with DTW assistance for 1 to 5 epochs.
-We ensure all experiments had same maximum numbers of training epochs and an identical early stopping strategy.
-Table 6 shows that the model’s predictive performance is poorest when DTW is not employed; moreover, no better results were obtained when DTW-assisted training exceeded one epoch.
-This finding confirms our expectation that, without DTW support during the initial stage, the model struggles to discover high-quality retrieval vectors.
-Concurrently, our main experiment has previously demonstrated that DTW is not inherently an optimal retrieval scheme geared towards forecasting accuracy.
-Consequently, prolonging the learning period under DTW guidance can actually restrict the expressive capacity of the forecasting model rather than improving it.
-Table 6: Effectiveness of Using DTW as Auxiliary Training.
+RATSF with embedding mae 1101.172
+mse 3343793.904
+decoding RATSF with encoder mae 1004.326
+4.3.2. Retrieval Embedding
+As previously mentioned, in our main experiments, we aimed to isolate the contribution of the encoder’s output to the retrieval performance.
+To do this, we utilized the encoder’s output $\mathbf{X}^{e}_{r}$ for the retrieval process and the embedding output $\mathbf{X}^{b}_{r}$ for decoding.
+In this section, we have completed the discussion of this part of the logic.
+Here, we will compare different schemes for retrieval and evaluate the effects of various decoding schemes.
+We argue that DTW is not the optimal retrieval method and demonstrate this through an experiment on the FHSV dataset, where we compare DTW, a separately trained two-layer MLP, and our current design choice: using the forecasting model’s encoder for retrieval.
+The upper half of Table 3 presents a comparison of the forecasting precision achieved with these different retrieval embedding strategies.
+The ’Baseline’ indicates the standard Transformer model without any retrieval mechanism.
+As the results indicate, the encoder output outperforms the MLP, likely due to the attention mechanism within the Transformer’s encoder, which is adept at identifying and refining the relationships among data points in sequences, thus capturing more accurate patterns.
+The lower half of Table 3 compares the decoding performance of RACA when using the Transformer’s embedding output and encoder output as the information source.
+The experimental results indicate that refining the retrieved sequence information with the Encoder can reduce the MAE by 18.38% to 1004.326, compared to the vanilla Transformer, and by 8.79% compared to using the embedding output.
+This substantiates the effectiveness of the RATSF architecture demonstrated in Figure 1.
+Table 4. Effectiveness of Using DTW as Auxiliary Training.
 dataset number of epochs metric
@@
 FHSV end to end mae 1302.054
+4.3.3. Effectiveness of Using DTW as Auxiliary Training
+In Section 3.3.2, we assert the effectiveness of employing DTW as a retrieval method during the cold start phase to facilitate the training of forecasting models for a single epoch.
+To substantiate the rationale for this strategy, we conducted a comparative experimental study: one baseline group without DTW-assisted training (referred to as ’end-to-end’), and several other groups, each trained with DTW assistance for 1 to 5 epochs.
+We ensured that all experiments had the same maximum number of training epochs and utilized an identical early stopping strategy.
+Table 4 demonstrates that the model’s predictive performance is the weakest when DTW is not utilized, as indicated by the highest MSE and MAE values in the last row of Table 4; furthermore, extending DTW-assisted training beyond one epoch did not lead to improved outcomes.
+This result aligns with our hypothesis that in the absence of DTW support during the initial training phase, the model has difficulty identifying high-quality retrieval vectors, which are crucial for accurate forecasting.
+However, as our primary experiments have shown, DTW is not inherently the best retrieval method for enhancing predictive accuracy.
+Therefore, extending the period of DTW-assisted training may hinder rather than enhance the model’s ability to express and forecast effectively.
+5. Conclusion
+In conclusion, RATSF stands out in the domain of univariant time-series forecasting by integrating a TSKB that meticulously slices and indexes historical data, leading to superior retrieval precision.
+This is further enhanced by the Encoder’s representations, which act as a sophisticated retrieval mechanism.
+Coupled with our innovative RACA mechanism for adeptly merging retrieved segments, we achieve heightened forecast accuracy.
+The adaptability of our method is highlighted by its compatibility with diverse Transformer variants and its extensive, successful application within Fliggy’s service volume forecasting scenarios.